OpenStudy (anonymous):

Double check my work?? Multiplying monomials (8x^2)^3 = 24x^6 (xyz^2) ^4 = x^5y^5z^8 d^3 times d^5 = d^8

OpenStudy (anonymous):

Show your work first..

OpenStudy (anonymous):

I just pretty much distribute for the first 2, and add 5 + 3 for the last one

OpenStudy (anonymous):

Last one is right.. You have to just add the exponents..

OpenStudy (anonymous):

There is little mistake in first one..

OpenStudy (anonymous):

I was taught when the exponent is outside the parenthesis, you multiply it to the inside

OpenStudy (anonymous):

Yes but you don't do like this: \(8^3 \ne 3 \times 8\) \(\large 8^3 = 8 \times 8 times 8\)

OpenStudy (anonymous):

\(\large 8^3 = 8 \times 8 \times 8\)

OpenStudy (anonymous):

Oh okay!